OpenStudy (anonymous):

Convert to a fraction in lowest terms. 15.4% I got this. Is this correct? 15.4% = 154/100 = 1 54/100 =1 27/50 The answer is 1 27/50

hero (hero):

It's 15.4/100, not 154/100

OpenStudy (anonymous):

= 1 54/100 This step is false.

OpenStudy (anonymous):

what you should have done is (15.4)/100

OpenStudy (anonymous):

Now, how can we remove the decimal point legally? (Hint: we're going to go to 154, but through a legal step in between)

OpenStudy (anonymous):

ok. i got it know thanks. I didn't out the decimal which made the entire problem wrong.

OpenStudy (anonymous):

Okie dokie. So you did do 154/1000 then?

OpenStudy (anonymous):

i meant I didn't put the decimal nor count the space's to reach the denominator.

OpenStudy (anonymous):

15.4% = 15.4/100 = 154/1000 =6. 49 Or 6 76/1000 =6.49 Or 6 16/250 The answer is 6 16/250

OpenStudy (anonymous):

On the step after 154/1000, that's not a legitimate thing. Your fraction is saying "154 divided by 1000", which equals (not surprisingly), .154. However, you misunderstood it to mean 1000/154 it seems. What you should do is go back to 154/1000. Since both are even, you should try simplifying by dividing both by 2. So, 77/500. To the best of my knowledge, this is as simple as you can make it. And if you test it by dividing 77/500, you still get .154.

OpenStudy (anonymous):

o Ok thanks for clarifying

OpenStudy (anonymous):

No problemo!